Hi,
I have 1/T = 0.00263
how do I find T? I don't remember the rule and can't find it anywhere!
thanks,
Kevin

Answer by jake_6233(74)

1/T = 0.00263 Remembering that if a/b=c/d then ad=bc
.00263T=1
T= 1/.00263 Also notice that you can flip both fractions as well